f(x)=1/2x Find f(0)
Answer by harpazo(655)

Replace very x you see with 0 and solve for f(0).
f(x)= 1/2x
f(0) = 1/2(0)
f(0) = 1/0
We cannot have a zero in the denominator of a fraction because we simply cannot divide by zero.
In conclusion, f(0) does not exist or it is undefined.
Understand?